10. Find the values of k for which the function f(x) = −4x 2 + (4k − 1)x − k 2 + 4 is negative for all values of x.


1
Expert's answer
2021-05-04T13:15:48-0400

The function "f(x)=-4x^2+(4k-1)x-k^2+4" is a parabola. Therefore, "f" does not intersect the abscissa axis, i.e. "-4x^2+(4k-1)x-k^2+4=0" has no solution.


Therefore, the discriminant must be negative:


"\Delta =b^2-4ac<0\Rightarrow (4k-1)^2-4\cdot (-4)\cdot (-k^2+4)<0"

"\Rightarrow (4k)^2-2\cdot 4k\cdot 1+1^2+16(-k^2+4)<0"

"\Rightarrow 16k^2-8k+1-16k^2+64<0\Rightarrow 65<8k"

"\Rightarrow k>\dfrac{65}{8}"

Therefore, the values of "k" that satisfy the condition are those that satisfy that "k>\dfrac{65}{8}" .